-3x + 8y = 48. Can someone please help me. Thanks!
Answer by tutorcecilia(2152)

-3x + 8y = 48 [Solve for y]
-3x+3x+8y=+3x+48
8y=3x+48
8/8y=3/8x+48/8
y=3/8x+6